The firm has $878,000 of sales and $913.000 of total assets. The firm is operating at 93 percent of capacity Capital Intensity Total Assets/Sales.) What is the capital intensity ratio at full capacity? 0.62 O b0.88 0.97 c. O d. 1.03 1.14
Answer #1

Sales at 93% capacity = $878000.

Therefore, sales at100% capacity would be = (878000/93)x100 = $944086.

Capital intensity ratios = total assets / sales = 913000/944086 = 0.97

Thus, correct option is (c) 0.97